{sigma x,sigma y}=0



1
Expert's answer
2021-04-22T07:27:06-0400

The Pauli matrices are

"\sigma_x = \begin{pmatrix}
 0 & 1\\
 1 & 0
\end{pmatrix}"

"\sigma_y = \begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix}"

"\sigma_z =\begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix}"

"\sigma_x \sigma _y = i \sigma_z"

"\sigma_y \sigma_x = - \sigma_x \sigma_y = -i \sigma_z"

The anticommutator of "\sigma_x" and "\sigma_y" is

"\{\sigma_x, \sigma_y\} = \sigma_x \sigma_y + \sigma_y \sigma_x = \sigma_x \sigma_y - \sigma_x \sigma_y =0"